Camphor is a white solid that melts at 179.5°C. It has been used to determine the molecular weights of organic compounds because of its unusually large freezing-point-depression constant (40°C/m), which allows ordinary thermometers to be used. The organic substance is dissolved in melted camphor, and then the melting point of the solution is determined.


1
Expert's answer
2022-03-04T16:43:04-0500

The freezing point depression (ΔTf) is related to the molality (m) and the freezing point depression constant (Kf) by ΔTf = K× m (for non-electrolyte compounds).

ΔTf = 179.8°C - 176.7°C = 3.1°C

Hence,

m = ΔTf / Kf = (3.1°C) / (40.00°C m–1) = 0.0775 m = 0.0775 mol/kg

m = 0.0775 mol/kg


Molality (m) = Moles of solute / Mass of solvent (kg),

Hence,

Moles of solute = Molality × Mass of solvent (kg) = 0.0775 mol/kg × 0.02201 kg = 0.001706 mol


Mass of solute = Moles of solute × Molar mass of solute

Hence,

Molar mass of solute = Mass of solute / Moles of solute

Molar mass of solute = (0.186 g) / (0.001706 mol) = 109.027 g/mol = 109 g/mol

Molar mass of solute = 109 g/mol


Answer: The molar mass of the unknown compound is 109 g/mol.